Team work 1 The choice of transportation mode The Carry- All luggage Company produces a line of luggage goods. The typical distribution plan is to produce a finished goods inventory located at the plant site carierhen shipped to company-owned field warehouses by way of common Rail is currently used to ship between East Coast plant and West Coast warehouse. The average transit time for rail shipments is 21 days. At each stocking point there is an average of 100 000 anits of luggage having an average production cost ofi(so USD per unit. Inventory holding costs are 30 %per year. There are 700 000 units sold per year out of the West Coast warehous Company is not sure that transportation by trains was the best decision Its obvious that acceleration of transportation and as result of that increasing shipments frequency must lead to proportional reduction of inventories on both sites of logistics chain. At the same time the fraction of yearly consumption which will be transit ( inventory in transit ) although decreases The company can use the following transport services Number of shipments per year Transport Rate Transit time ($ / unit ) (days ) 21 14 Rail Piggyback Truck 0.10 0.15 0.20 1.40 10 20 20 40 ir 2 Assignment You are asked to help company to make right decision in transportation choice. Ordering costs are assumed to be negligible.

please help asap!
Answer #1

Ans) Explanations for the formula stated in the below chart

a)Transport costs: If R denotes unit transport rate and D denotes annual demand, then RD gives an estimate of the annual transport cost.

b)In-transit inventory: Each unit, valued at $C, spends T days in transit. If I is annual holding rate, each item costs ICT/365 in holding charges during transport time. Since D is annual demand, total in-transit inventory cost equals [ICT/365]*D.

c)Letting Q = shipment size, and assuming production occurs instantaneously at the plant, average annual inventory cost at the plant equals ICQ/2.(for rail Q/2=100 both at the plant and at the warehouse)

d) Letting C ’ = C + R, i.e., product value at the field warehouse, then average inventory cost at the field warehouse equals IC’ ‘Q/2, assuming constant lead time (All of these inventory cost estimates assume constant and deterministic demand rate)

Cost types-

i)For transportation; (Formula- R*D)

Rail- (0.1)(700.000) =70.000 Piggyback- (0.15)(700.000) =105.000 Truck- (0.2)(700.000) =140.000 Air- (1.4)(700.000) =980.000

ii)For In-transit Inventory; (Formula- ICDT/365)

Rail- ((0.3)(30) X(700 00)X(21))/365 =363.465  Piggyback- ((0.3)(30) X(700.000) X(14))/365 =241.644  Truck- ((0.3)(30) X(700.000)X(5))/36 5 =86.301 Air- ((0.3)(30) X(700.000)X(2))/3 65 =34.521

iii)For Plant Inventory; (Formula- ICQ/2)

Rail- ((0.3)(30) X(100.000)) =900.000  Piggyback- ((0.3)(30) X(50.000)(0.93) =418.500   Truck-  ((0.3)(30) (50.000) (0.84)) =378.000  Air- (0.3)(30) X(25.000) X(0.8) =182.250

iv)For Warehouse Inventory; (Formula-IC’Q/2)

Rail- ((0.3)(30.1) X(100.000)) =903.000  Piggyback- ((0.3)(30.15) X(50.000)(0.93) =420.593   Truck- ((0.3)(30.2) (50.000 X(0.84)) =380.520   Air- ((0.3)(30.4) X(25000) X(0.8)) =190.755

Therefore Total Cost- (adding all cost types for the particular transport)

Rail- 2236.465

Piggyback- 1185.737

Truck- 984.821

Air- 1387.526

Conclusion; Transportation cost trough Truck will cost the lowest for the company.